# An Agentic Operationalization of DISARM for FIMI Investigation on Social Media

### v1

@@ III-D1 Exploratory data analysis (EDA)
The workflow begins with its first iteration by ingesting raw social media data into a SQL-queryable format. The first iteration performs an EDA that captures the dataset’s statistics–sums, unique accounts, temporal range, and platform-specific metadata–without conducting anomaly detection. This deferred approach ensures that subsequent findings are attributed to specific DISARM techniques rather than generic anomalies.
@@ III-D2 TTPs sampling and interpretation
Beyond the first iteration of EDA, the LLM agent samples a DISARM technique using a multi-arm bandit [7] explore-exploit strategy:
  • Explore mode (early iterations): Prioritize untested techniques to maximize coverage of the DISARM taxonomy.
  • Exploit mode (later iterations): Deep-dive into techniques showing promising signals from prior iterations.
@@ IV-C1 Chinese campaign on X
The key finding is the detection of a suspicious burst in account creation dates that, after evaluation, we corroborate as being related to the creation of IO accounts. This pattern was previously observed in manual analysis of the full dataset. Most accounts created in August 2018 are IO accounts (precision = 0.97), demonstrating the agent-picked algorithm’s ability to surface temporally clustered coordinated account creation.
Additional signals refining this pattern–such as empty profile descriptions, campaign-related keywords, and interaction patterns–do not substantially improve detection given the already high precision of the parent signal. However, they provide useful characterizations of the subgroups of the IO account, including the characteristics of activity and content.
The main signal absent from agent-generated evidence is temporal synchronization. For example, two users repeatedly retweet the same posts within a short time window (e.g., 10 minutes) over multiple days. This type of coordination is present in the dataset for retweets, replies, and especially mentions, but the current pipeline does not capture the time dimension.
@@ IV-C2 Russian campaign on Telegram
It also captured several auxiliary metrics that aligned with patterns previously observed by human analysts. These included, among others, duplicated comments, shorter average comment length, and a lower mean number of reactions received by bot-generated comments. While such features proved useful in corroborating hypotheses regarding the artificial nature of borderline or ambiguous accounts (“gray-zone”), they were insufficiently robust to serve as primary indicators for bot detection when considered in isolation.
At the same time, certain indicators employed during human-led identification did not emerge among the algorithmically selected features. Notably, these included measures such as consecutive days of sustained activity and the repeated dissemination of identical links to AI-generated video content produced as part of the information operation. This discrepancy highlights both the complementary strengths and the current limitations of automated approaches compared to expert-driven qualitative assessment.
@@ IV-D Cost & Efficiency
The workflow was executed using Claude Code. When running on Claude Opus 4.5, the process was deliberately split into segmented runs to prevent short-cutting of iterative steps. In total, 15 iterations were performed, including an initial iteration dedicated to exploratory data analysis. The total API runtime was approximately 35 minutes, with a total cost of USD 11.4.
@@ V Discussion
At this stage, output of the automated analysis should be understood as a structured set of techniques and patterns that can support bot identification rather than as a fully autonomous detection system. CIB detection is typically based on a rule-based decision framework in which heterogeneous signals - both behavioral and content-based - carry different weights and are combined to classify individual accounts. In this study, the automated approach produced a consolidated table of TTPs and their operationalization within this specific dataset. However, the absence of an explicit decision tree or a weighted set of rules underscores the continued need for human judgment in the final classification process.
@@
Not all proposed metrics proved equally effective in this dataset. Several thresholds required additional contextual constraints to be informative. For instance, message duplication is a well-established signal of coordinated activity, yet its practical use depends on factors such as message length, as short or emoji-based comments may be frequently repeated by organic users. Similarly, the message length itself emerged as a potentially relevant signal, but, as noted in the Results section, remains insufficient as a standalone indicator.
@@
A focus on purely behavioral patterns brings about both advantages and limitations. Its primary strength lies in its relative universality and reduced dependence on the linguistic or political context. At the same time, this approach limits analytical resolution in several important respects. Behavioral signals alone do not allow researchers to distinguish between multiple bot farms operating simultaneously, to separate coordinated groups acting in opposition to one another, or to differentiate between bots and trolls (the latter referring to human-operated accounts engaged in coordinated or semi-coordinated manipulation). Although generative AI is blurring the boundary between these categories, in this dataset they remain associated with distinct content characteristics–automatically generated versus human-authored–which is a distinction that remains operationally relevant for designing effective countermeasures or statements on inauthenticity.
@@
These limitations point to the potential value of incorporating a temporal dimension into automated detection. Longitudinal CIB operations evolve over time, as do the patterns through which they manifest. In this case, the same bot infrastructure has been observed operating in other contexts, including Russian-occupied territories of Ukraine [2], with partially overlapping accounts. During the past year, this operation appears to have shifted from highly repetitive commenting strategies to more diverse AI-generated messages. Capturing such shifts through temporal modeling could improve long-term detection and make adaptation itself a detectable signal of coordinated influence.
@@
Despite these constraints, human evaluation suggests that the techniques and patterns identified by agentic AI remain valuable as a research aid. Rather than replacing expert analysis, they function as a structured guide that highlights relevant signals and narrows the analytical focus, supporting a more consistent and transparent human-led assessment of coordinated inauthentic behavior.
@@ V-B Limitations and future work
While the agentic system excels at recognition of behavioral patterns, it is currently limited in capturing temporal synchronicity and nuanced linguistic shifts. Future work should focus on modeling temporal evolution to detect shifting campaign tactics and integrating multi-modal content analysis to distinguish between automated bots and human-operated trolls. Additionally, to further confirm the generalization of the method, the pipeline should be evaluated against a broader range of social media platforms and image-based influence campaigns.
@@ VI Conclusion
This paper presents a multi-agent pipeline to operationalize the DISARM framework for augmenting FIMI investigation on social media. We demonstrated a novel approach where autonomous AI agents collaboratively detect manipulative behaviors and map them to standardized TTPs.
Our evaluation on two real-world datasets yielded a 50% technique pass rate across 14 autonomous iterations. In particular, the system successfully identified more than 30 previously undetected bot accounts in the Telegram dataset, proving its ability to surface actionable intelligence beyond initial human benchmarks. These results confirm that autonomous agents can function as a cost-effective force multiplier–compressing the investigation timeline while ensuring outputs remain doctrinally aligned for interoperability. While human oversight remains essential for assessing intent, this work establishes a verifiable baseline for automated FIMI investigation in high-tempo hybrid defense environments.

### v2

@@ III-D1 Exploratory data analysis (EDA)
The workflow begins with its first iteration by ingesting raw social media data into a SQL-queryable format. The first iteration performs an EDA that captures the dataset’s statistics–sums, unique accounts, temporal range, and platform-specific metadata–without conducting anomaly detection. This deferred approach avoids premature attribution, allowing DISARM alignment to be established only in the later technique-guided iterations.
@@ III-D2 TTPs sampling and interpretation
Beyond the first iteration of EDA, the LLM agent samples a DISARM technique using an LLM-guided multi-arm bandit [7] explore-exploit strategy:
  • Explore: Prioritize untested techniques to maximize coverage of DISARM taxonomies.
  • Exploit: Deep-dive into techniques showing promising signals from prior iterations.
  • Strategy adaptation: The explore-exploit balance is dynamically adjusted based on cumulative hunt outcomes—shifting toward explore after repeated failures, and toward exploit after discovering strong signals.
@@ IV-C1 Chinese campaign on X
The key finding is the detection of a suspicious burst in account creation dates that, after evaluation, we corroborate as being related to the creation of IO accounts. This pattern was previously observed in manual analyses of the full dataset. Most accounts created in August 2018 are IO accounts (precision = 0.97), demonstrating the agent-picked algorithm’s ability to surface temporally clustered, coordinated account creation.
Additional signals refining this pattern—such as empty profile descriptions, campaign-related keywords, and interaction patterns—do not substantially improve detection, given the already high precision of the parent signal. However, they provide useful characterizations of subgroups of the IO accounts, including the characteristics of their activity and content.
The main signal absent from agent-generated evidence is temporal synchronization. For example, two users repeatedly retweet the same posts within a short time window (e.g., 10 minutes) over multiple days. This type of coordination is present in the dataset for retweets, replies, and especially mentions, but the current pipeline does not capture the temporal dimension.
@@ IV-C2 Russian campaign on Telegram
It also captured several auxiliary metrics that aligned with patterns previously observed by human analysts. These included, among others, duplicated comments, shorter average comment length, and a lower mean number of reactions received by bot-generated comments. While such features proved useful in corroborating hypotheses regarding the artificial nature of borderline or ambiguous (“gray-zone”) accounts, they were insufficiently robust to serve as primary indicators for bot detection when considered in isolation.
At the same time, certain indicators employed during human-led identification did not emerge among the algorithmically selected features. Notably, these included measures such as consecutive days of sustained activity and the repeated dissemination of identical links to AI-generated video content produced as part of the information operation. This discrepancy highlights both the complementary strengths and the current limitations of automated approaches compared with expert-driven qualitative assessment.
@@ IV-D Cost & Efficiency
The workflow was executed using Claude Code. When running on Claude Opus 4.5, the process was deliberately split into multiple runs to prevent shortcutting iterative steps. In total, 15 iterations were performed, including an initial iteration dedicated to exploratory data analysis. The total API runtime was approximately 35 minutes, with a total cost of USD 11.40.
@@ V Discussion
At this stage, the output of the automated analysis should be understood as a structured set of techniques and patterns that can support bot identification, rather than as a fully autonomous detection system. CIB detection is typically based on a rule-based decision framework in which heterogeneous signals—both behavioral and content-based—carry different weights and are combined to classify individual accounts. In this study, the automated approach produced a consolidated table of TTPs and their operationalization within this specific dataset. However, the absence of an explicit decision tree or a weighted set of rules underscores the continued need for human judgment in the final classification process.
@@
Not all proposed metrics proved equally effective in this dataset. Several thresholds required additional contextual constraints to be informative. For instance, message duplication is a well-established signal of coordinated activity, yet its practical use depends on factors such as message length, as short or emoji-based comments may be frequently repeated by organic users. Similarly, message length itself emerged as a potentially relevant signal but, as noted in the Results section, remains insufficient as a standalone indicator.
@@
A focus on purely behavioral patterns brings both advantages and limitations. Its primary strength lies in its relative universality and reduced dependence on linguistic or political context. At the same time, this approach limits analytical resolution in several important respects. Behavioral signals alone do not allow researchers to distinguish between multiple bot farms operating simultaneously, to separate coordinated groups acting in opposition to one another, or to differentiate between bots and trolls (the latter referring to human-operated accounts engaged in coordinated or semi-coordinated manipulation). Although generative AI is blurring the boundary between these categories, in this dataset they remain associated with distinct content characteristics—automatically generated versus human-authored—which is a distinction that remains operationally relevant for designing effective countermeasures or statements on inauthenticity.
@@
These limitations point to the potential value of incorporating a temporal dimension into automated detection. Longitudinal CIB operations evolve over time, as do the patterns through which they manifest. In this case, the same bot infrastructure has been observed operating in other contexts, including Russian-occupied territories of Ukraine [2], with partially overlapping accounts. Over the past year, this operation appears to have shifted from highly repetitive commenting strategies to more diverse AI-generated messages. Capturing such shifts through temporal modeling could improve long-term detection and make adaptation itself a detectable signal of coordinated influence.
@@
Despite these constraints, human evaluation suggests that the techniques and patterns identified by agentic AI remain valuable as research aids. Rather than replacing expert analysis, they function as structured guides that highlight relevant signals and narrow the analytical focus, supporting a more consistent and transparent human‑led assessment of coordinated inauthentic behavior.
@@ V-B Limitations and future work
While the agentic system excels at recognizing behavioral patterns, it is currently limited in capturing temporal synchronicity and nuanced linguistic shifts. Future work should focus on modeling temporal evolution to detect shifting campaign tactics and on integrating multimodal content analysis to distinguish between automated bots and human-operated trolls. Additionally, to further increase the method’s generalizability, the pipeline should be evaluated across a broader range of social media platforms and image-based influence campaigns.
@@ VI Conclusion
This paper presents an agentic pipeline to operationalize the DISARM framework for augmenting FIMI investigations on social media. We demonstrate a novel approach in which autonomous AI agents collaboratively detect manipulative behaviors and map them to standardized TTPs.
Our evaluation on two real-world datasets yielded a 50% technique pass rate across 14 autonomous iterations. In particular, the system successfully identified more than 30 previously undetected bot accounts in the Telegram dataset, demonstrating its ability to surface actionable intelligence beyond initial human benchmarks. These results confirm that autonomous agents can function as a cost-effective force multiplier, compressing the investigation timeline while ensuring outputs remain doctrinally aligned for interoperability. While human oversight remains essential for assessing intent, this work establishes a verifiable baseline for automated FIMI investigations in high-tempo hybrid defense environments.